OpenStudy (anonymous):

Help** A land developer sold 3 3/4 acres of his 18 3/5 acres. How many acres does he have left?

OpenStudy (anonymous):

18 and 3/5 minus 3 and 3/4 is equal to?

OpenStudy (anonymous):

Convert these into improper fractions and subtract by converting the bottom denominators into the same denominator

OpenStudy (anonymous):

3 and 3/4=15/4 18 and 3/5=93/5

OpenStudy (anonymous):

A common denominator would be 20. Since 20 is divisible by both 4 and 5.

OpenStudy (anonymous):

So 15/4=75/20 and 93/5=372/20.

OpenStudy (anonymous):

(372-75)/20 is equal to?

OpenStudy (anonymous):

297/20!

OpenStudy (anonymous):

Thank you for explaining that!